# ForceGrip: Data-Free Curriculum Learning for Realistic Grip Force Control in VR Hand Manipulation

### v1

@@ Abstract
Realistic Hand manipulation is a key component of immersive virtual reality (VR), yet existing methods often rely on a kinematic approach or motion-capture datasets that omit crucial physical attributes such as contact forces and finger torques. Consequently, these approaches prioritize tight, one-size-fits-all grips rather than reflecting users’ intended force levels. We present ForceGrip, a deep learning agent that synthesizes realistic hand manipulation motions, faithfully reflecting the user’s grip force intention. Instead of mimicking predefined motion datasets, ForceGrip uses generated training scenarios-randomizing object shapes, wrist movements, and trigger input flows—to challenge the agent with a broad spectrum of physical interactions. To effectively learn from these complex tasks, we employ a three-phase curriculum learning framework comprising Finger Positioning, Intention Adaptation, and Dynamic Stabilization. This progressive strategy ensures stable hand-object contact, adaptive force control based on user inputs, and robust handling under dynamic conditions. Additionally, a proximity reward function enhances natural finger motions and accelerates training convergence. Quantitative and qualitative evaluations reveal ForceGrip’s superior force controllability and plausibility compared to state-of-the-art methods.
Figure 1: Overview of our training pipeline. Rather than relying on motion-capture data, we generate diverse scenarios of varying object shapes, user inputs, and wrist motions; a three-phase curriculum progressively increases task complexity.
@@
Keywords: Virtual reality, Hand manipulation, Interaction, Physics-based animation, Curriculum learning
1 Introduction
@@
Immersive hand interaction is vital for compelling virtual reality (VR) experiences, enabling users to naturally grasp and manipulate objects. Although controller-based interfaces have become prevalent due to their accessibility, usability and responsiveness Han et al. (2023), most existing solutions overlook the nuanced control of grip force, focusing instead on reproducing visually plausible hand motions. This gap undermines a key dimension of realistic interaction: in physical environments, users continually fine-tune their grip force to accommodate differences in object weight, fragility, and friction, allowing objects to slip under varying force levels or remain firmly secured when needed.
@@
Translating these nuances into VR poses considerable challenges, as conventional datasets and motion-capture techniques rarely include detailed physical attributes such as contact forces or finger torques. Consequently, it is crucial to provide diverse scenarios that involve various physical demands and ensure that these complex scenarios can be effectively learned.
@@
  1. I.
    Dataset-free training. ForceGrip synthesizes visually and physically realistic hand manipulations without relying on motion-capture data.
  2. II.
@@
2 Related Works
2.1 VR Interfaces for Physical Interaction Modeling
Two principal approaches exist for hand interactions in VR: hand-tracking and VR controller interfaces. Hand-tracking interfaces rely on sensors, such as VR gloves or vision-based sensors, offering high embodiment by mirroring the user’s actual hand movements Voigt-Antons et al. (2020). However, they often use penetration-based models for grip force Höll et al. (2018); Quan et al. (2020), which require careful per-finger calibration and can undermine usability and confidence Voigt-Antons et al. (2020); Kangas et al. (2022). Furthermore, task performance depends on the sensing hardware quality Viola et al. (2022), restricting the general user’s accessibility.
In contrast, VR controller interfaces treat the controller’s pose as the wrist proxy, while finger motions are either explicitly generated or inferred. A basic technique is the attachment-based method Oprea et al. (2019); Shi et al. (2022), which triggers fixed finger-closing motions and attaches objects upon collision. Although intuitive to use, it lacks motion diversity and physical realism. VR-HandNet Han et al. (2023) addresses some of these issues by inferring joint torques from virtual sensor data and comparing them to motion-capture datasets. However, because these datasets often lack physical parameters—such as contact forces and finger torques—the system’s ability to provide fine-grained grip force control remains constrained.
@@ 2.1 VR Interfaces for Physical Interaction Modeling
Table 1: Structured comparison of existing VR hand manipulation methods by presenting whether they are controller-based, involve physics simulation, employ learning-based strategies, or require a reference motion dataset.
@@ X O X O
2.2 Physics-based Animation Training
Physics-based animation calculates joint torques via a physics engine to account for collisions, balance, and other constraints, enabling systems to adapt dynamically to environmental changes and external forces. Popular engines Todorov et al. (2012); Coumans and Bai (2016); Liang et al. (2018); Makoviychuk et al. (2021) often lack differentiability or smooth integration with deep learning. Although differentiable engines exist Hu et al. (2019); Freeman et al. (2021), they see limited use due to computational overhead and restricted physics features. Consequently, reinforcement learning (RL) Sutton and Barto (1999) serves as the dominant method, as it trains agents via reward signals without requiring differentiability.
@@ 2.2 Physics-based Animation Training
In animation reproduction research, the objective is to replicate target motions within a physically simulated environment for enhanced realism. Techniques such as imitation reward Peng et al. (2017, 2018); Park et al. (2019); Won and Lee (2019); Lee et al. (2021b) and generative adversarial imitation learning (GAIL) Ho and Ermon (2016) are often adopted to align simulated poses with reference motion data Luo et al. (2020); Peng et al. (2021, 2022); Hassan et al. (2023). This approach preserves natural motion quality but can be limited by the coverage of available datasets.
In task completion research, agents learn specific tasks such as locomotion Heess et al. (2017); Klipfel et al. (2023), object manipulation Kumar and Todorov (2015); Andrychowicz et al. (2020); Christen et al. (2022); Zhang et al. (2024), or specialized actions like door opening Rajeswaran et al. (2017) or chopstick handling Yang et al. (2022).
In user control research, real-time user inputs dynamically reshape the agent’s objectives, integrating animation reproduction for smooth, natural responses. Examples include determining walking directions Bergamin et al. (2019); Wang et al. (2020); Won et al. (2022), generating diverse skill animations Lee et al. (2021a), and controlling animations through natural language Juravsky et al. (2022); Cui et al. (2024). For hand-specific control, VR controllers can provide user signals used for finger motion simulation Han et al. (2023).
@@
2.3 Curriculum Learning
Curriculum learning Bengio et al. (2009) organizes training tasks in increasing order of difficulty, a strategy shown to accelerate convergence and improve performance in deep learning Weinshall et al. (2018); Baker et al. (2019); Hacohen and Weinshall (2019).
Within physics-based animation, this approach helps agents learn complex skills more efficiently. For example, Won et al. (2021) first trained basic skills and then introducing adversarial environments. Similarly, Liu et al. (2022); Luo et al. (2023); Haarnoja et al. (2024) guides agents to learn increasingly diverse and challenging skills. Although automated task sequence generation has been explored Graves et al. (2017); Florensa et al. (2018); Silva and Costa (2018); Narvekar and Stone (2018), no universal standard exists for defining task difficulty Narvekar et al. (2020). Instead, manual structuring remains highly effective Justesen et al. (2018); Cobbe et al. (2019); Wang et al. (2019).
@@ 2.3 Curriculum Learning
3 Overview
Figure 1 depicts our training pipeline. We employ Unity with NVIDIA PhysX for real-time physics, updating at 120 Hz while the learning agent acts at 30 Hz. Each time step $t$, the agent outputs torques for finger joints, which a proportional-derivative (PD) controller Tan et al. (2011) translates into hand poses and drives object interactions at time $t+1$. During training, the wrist follows pre-set motion paths, whereas in actual usage it follows VR controller movements.
Instead of relying on motion-capture data, we generate diverse scenarios encompassing varied object shapes, user inputs, and wrist motions to prevent overfitting and enhance realism. However, this variability increases learning complexity. To address this, our curriculum strategy partitions training into simpler grip tasks, followed by progressively more dynamic conditions.
To train the agent, we leverage RL. We use proximal policy optimization (PPO) Schulman et al. (2017), running 576 concurrent agents with different seeds to maximize training diversity. Detailed hyperparameter settings are given in the supplementary materials. Additionally, we employ early termination Peng et al. (2018) to stop an episode if an object moves more than 10 cm away from the wrist, avoiding wasted computation on irrecoverable states. This mechanism increases efficiency and prevents spurious rewards, ultimately leading to a more stable policy that can handle the intricate demands of VR grip force control.
Figure 2: (a) Hand model from Meta XR Core SDK. (b) Objects: sphere, cube, and cylinder. (c) Sensors: GVS, LVS, closest and normal vector.
3.1 Hand Model and Objects
We utilize the Meta XR Core SDK hand model, which has 17 finger joints (23 total DoFs) and 19 capsule colliders spread across the fingers and palm (Figure 2(a)). To learn various grip conditions, we use three primitive object shapes—sphere, cube, and cylinder (Figure 2(b))—each with a uniform mass of 0.01 kg. Rather than ensuring all objects can always be lifted, our objective is to reflect user-intended force; insufficient force causes objects to slip. However, since dropping an object triggers early termination and thus biases the agent to apply overly strong grips, we mitigate mass effects by keeping these objects lightweight.
3.2 State and Action
@@
Hand model information $H=\{p,q,v,w,\alpha\}$ includes joint positions $p\in\mathbb{R}^{66}$, joint DoF angles $q\in\mathbb{R}^{23}$, joint linear velocities $v\in\mathbb{R}^{51}$, joint DoF velocities $w\in\mathbb{R}^{23}$, and joint DoF accelerations $\alpha\in\mathbb{R}^{23}$. All positional attributes are expressed in meters and shifted into a wrist-centric coordinate system, while angular values use radians to standardize feature ranges.
Object information $O=\{o_{g},v_{gvs},v_{lvs},c,n\}$ includes the gravity vector $o_{g}\in\mathbb{R}^{3}$ in wrist coordinates, a global voxel sensor $v_{gvs}\in\mathbb{R}^{450}$ attached to the wrist, employing 2 cm cells, local voxel sensors $v_{lvs}\in\mathbb{R}^{2160}$ attached to each collider, employing 0.5 cm cells, closest point vectors $c\in\mathbb{R}^{69}$ from relevant joints to the object’s surface, and corresponding surface normals $n\in\mathbb{R}^{69}$. Inspired by previous work Zhang et al. (2021); Han et al. (2023), this dual-resolution sensing strategy provides a balance between detail and generality, minimizing overfitting to a specific object shape. Object information is illustrated in Figure 2.
@@ 3.3 Reward function
$r_{p}=\sum_{j}{w^{j}\exp[-0.07||c_{t}^{j}||^{2}]}$ (3)
where $c_{t}^{j}$ is the distance vector from the $j$-th joint to the object surface, and $w^{j}$ is a weight emphasizing fingertip-first contact. We assign higher weights (0.0625) to the five end-effectors and their associated joints, and lower weights (0.03125) elsewhere.
@@
$r=r_{f}+r_{p}$ (4)
3.4 Model Architecture
At each timestep, the agent processes the state $s$ by splitting it into logically distinct components: $p,q,v,w,\alpha,o_{g},v_{gvs},v_{lvs},\allowbreak c,n,f,u,p$, embedding each subset with specialized encoder layers and concatenating the results into a 580-dimensional vector. This vector then passes through 10 fully connected layers (580 neurons each, GeLU Hendrycks and Gimpel (2016), producing 23 action values that control joint torques. Additional architectural details, including encoder configurations and hyperparameters, appear in the supplementary materials.
4 Training
4.1 Training scenario generation
@@
4.2 Curriculum learning
Because these random, high-complexity scenarios pose significant training challenges, we devise a three-phase curriculum: Finger Positioning, Intention Adaptation, and Dynamic Stabilization. Each phase builds on the skills developed in the preceding phase, progressively training the agent to handle increasingly dynamic and realistic conditions.
4.2.1 Phase 1: Finger Positioning
This phase trains the agent to establish accurate finger contact. To minimize distractions, only object randomization is active; trigger variation and wrist movement are disabled. The wrist remains fixed, and user triggers are set to constant random values within [0,1]. This setup accelerates the agent’s mastery of foundational behaviors such as collision handling and initial grip formation.
4.2.2 Phase 2: Intention Adaptation
@@
4.2.3 Phase 3: Dynamic Stabilization

### v3

# ForceGrip: Reference-Free Curriculum Learning for Realistic Grip Force Control in VR Hand Manipulation

@@ Abstract.
Realistic Hand manipulation is a key component of immersive virtual reality (VR), yet existing methods often rely on a kinematic approach or motion-capture datasets that omit crucial physical attributes such as contact forces and finger torques. Consequently, these approaches prioritize tight, one-size-fits-all grips rather than reflecting users’ intended force levels. We present ForceGrip, a deep learning agent that synthesizes realistic hand manipulation motions, faithfully reflecting the user’s grip force intention. Instead of mimicking predefined motion datasets, ForceGrip uses generated training scenarios—randomizing object shapes, wrist movements, and trigger input flows—to challenge the agent with a broad spectrum of physical interactions. To effectively learn from these complex tasks, we employ a three-phase curriculum learning framework comprising Finger Positioning, Intention Adaptation, and Dynamic Stabilization. This progressive strategy ensures stable hand-object contact, adaptive force control based on user inputs, and robust handling under dynamic conditions. Additionally, a proximity reward function enhances natural finger motions and accelerates training convergence. Quantitative and qualitative evaluations reveal ForceGrip’s superior force controllability and plausibility compared to state-of-the-art methods. Demo videos are available as supplementary material and the code is provided at https://han-dongheun.github.io/ForceGrip.
@@
Virtual reality, Hand manipulation, Interaction, Physics-based animation, Curriculum learning
journalyear: 2025copyright: ccconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers; August 10–14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers (SIGGRAPH Conference Papers ’25), August 10–14, 2025, Vancouver, BC, Canadadoi: 10.1145/3721238.3730738isbn: 979-8-4007-1540-2/2025/08ccs: Computing methodologies Virtual realityccs: Computing methodologies Physical simulation
Figure 1. ForceGrip is a deep learning agent that synthesizes realistic hand manipulation motions and faithfully reflects the user’s intended grip force.
[…]
A visual summary of the ForceGrip agent showing a hand manipulating an object, representing the agent’s ability to reflect user grip force.
1. Introduction
@@
Immersive hand interaction is vital for compelling virtual reality (VR) experiences, enabling users to naturally grasp and manipulate objects. Although controller-based interfaces —without relying on hand tracking— have become prevalent due to their accessibility, usability and responsiveness (Han et al., 2023), most existing solutions overlook the nuanced control of grip force. Instead, they emphasize visually plausible hand motions that rely on simple, binary trigger inputs. This gap undermines a key dimension of realistic interaction: in physical environments (Holonautic, 2021), users continually fine-tune their grip force to accommodate differences in object weight, fragility, and friction, allowing objects to slip under varying force levels or remain firmly secured when needed.
@@
Translating these nuances into VR poses considerable challenges, as conventional datasets and motion-capture techniques (Kim et al., 2024) rarely include detailed physical attributes such as contact forces or finger torques. Consequently, it is crucial to provide diverse scenarios that involve various physical demands and ensure that these complex scenarios can be effectively learned.
@@
  1. (1)
    Reference-free training. ForceGrip synthesizes visually and physically realistic hand manipulations without relying on motion-capture data.
  2. (2)
@@
2. Related Works
2.1. VR Interfaces for Physical Interaction Modeling
Two principal approaches exist for hand interactions in VR: hand-tracking and VR controller interfaces. Hand-tracking interfaces rely on sensors, such as VR gloves or vision-based sensors, offering high embodiment by mirroring the user’s actual hand movements (Voigt-Antons et al., 2020). However, they often use penetration-based models for grip force (Höll et al., 2018; Quan et al., 2020), which require careful per-finger calibration and can undermine usability and confidence (Voigt-Antons et al., 2020; Kangas et al., 2022). Furthermore, task performance depends on the sensing hardware quality (Viola et al., 2022), restricting the general user’s accessibility.
In contrast, VR controller interfaces treat the controller’s pose as the wrist proxy, with finger motions either predefined or inferred. A basic approach is the attachment-based method (Oprea et al., 2019; Shi et al., 2022), which closes fingers in fixed animations and attaches objects on contact. While intuitive, it lacks motion diversity and physical realism. VR-HandNet (Han et al., 2023) addresses this by inferring joint torques from motion-capture data, but the absence of physical attributes in the dataset limits precise grip force control. Meanwhile, physics-aware controllers (Lee et al., 2019; Gonzalez-Franco et al., 2020; Sinclair et al., 2019) and avatars (Tao et al., 2023) have been explored to address the limitations of standard VR controllers.
@@ 2.1. VR Interfaces for Physical Interaction Modeling
Table 1. Structured comparison of existing VR hand manipulation methods by presenting whether they are controller-based, involve physics simulation, employ learning-based strategies, or require a reference motion dataset.
@@ X O X O
[…]
A comparison table listing different VR hand manipulation methods. Rows represent different methods: Penetration, Attachment, VR-HandNet, and ForceGrip. Columns indicate whether each method is controller-based, physics-based, learning-based, and reference-free. Checkmarks (O) and crosses (X) show support for each feature. ForceGrip supports all four.
2.2. Physics-based Animation Training
Physics-based animation calculates joint torques via a physics engine to account for collisions, balance, and other constraints, enabling systems to adapt dynamically to environmental changes and external forces. Popular engines (Todorov et al., 2012; Coumans and Bai, 2016; Liang et al., 2018; Makoviychuk et al., 2021) often lack differentiability or smooth integration with deep learning. Although differentiable engines exist (Hu et al., 2019; Freeman et al., 2021), they see limited use due to computational overhead and restricted physics features. Consequently, reinforcement learning (RL) (Sutton and Barto, 1999) serves as the dominant method, as it trains agents via reward signals without requiring differentiability.
@@ 2.2. Physics-based Animation Training
In animation reproduction research, the objective is to replicate target motions within a physically simulated environment for enhanced realism. Techniques such as imitation reward (Peng et al., 2017, 2018; Park et al., 2019; Won and Lee, 2019; Lee et al., 2021a) and generative adversarial imitation learning (GAIL) (Ho and Ermon, 2016) are often adopted to align simulated poses with reference motion data (Luo et al., 2020; Peng et al., 2021, 2022; Hassan et al., 2023). This approach preserves natural motion quality but can be limited by the coverage of available datasets.
In task completion research, agents learn specific tasks such as locomotion (Heess et al., 2017; Klipfel et al., 2023), object manipulation (Kumar and Todorov, 2015; Andrychowicz et al., 2020; Christen et al., 2022; Zhang et al., 2024), or specialized actions like door opening (Rajeswaran et al., 2017) or chopstick handling (Yang et al., 2022).
In user control research, real-time user inputs dynamically reshape the agent’s objectives, integrating animation reproduction for smooth, natural responses. Examples include determining walking directions (Bergamin et al., 2019; Wang et al., 2020; Won et al., 2022), generating diverse skill animations (Lee et al., 2021b), and controlling animations through natural language (Juravsky et al., 2022; Cui et al., 2024). For hand-specific control, VR controllers can provide user signals used for finger motion simulation (Han et al., 2023).
@@
2.3. Curriculum Learning
Curriculum learning (Bengio et al., 2009) organizes training tasks in increasing order of difficulty, a strategy shown to accelerate convergence and improve performance in deep learning (Weinshall et al., 2018; Baker et al., 2019; Hacohen and Weinshall, 2019).
Within physics-based animation, this approach helps agents learn complex skills more efficiently. For example, (Won et al., 2021) first trained basic skills and then introducing adversarial environments. Similarly, (Liu et al., 2022; Luo et al., 2023; Zhang et al., 2023; Haarnoja et al., 2024) guide agents to learn increasingly diverse and challenging skills. Although automated task sequence generation has been explored (Graves et al., 2017; Florensa et al., 2018; Silva and Costa, 2018; Narvekar and Stone, 2018), no universal standard exists for defining task difficulty (Narvekar et al., 2020). Instead, manual structuring remains highly effective (Justesen et al., 2018; Cobbe et al., 2019; Wang et al., 2019).
@@ 2.3. Curriculum Learning
3. Overview
Figure 1 depicts our training pipeline. We employ Unity with NVIDIA PhysX for real-time physics, updating at 120 Hz while the learning agent acts at 30 Hz. Each time step $t$, the agent outputs torques for finger joints, which a proportional-derivative (PD) controller (Tan et al., 2011) translates into hand poses and drives object interactions at time $t+1$. During training, the wrist follows pre-set motion paths, whereas in actual usage it follows VR controller movements.
Instead of relying on motion-capture data, we generate diverse scenarios encompassing varied object shapes, user inputs, and wrist motions. However, this variability increases learning complexity. To address this, our curriculum strategy partitions training into simpler tasks, followed by progressively more dynamic conditions.
To train the agent, we leverage RL. We use proximal policy optimization (PPO) (Schulman et al., 2017), running 576 concurrent agents with different seeds. Detailed hyperparameter settings are given in the supplementary materials. We employ early termination (Peng et al., 2018) to stop an episode if an object moves more than 10 cm away from the wrist, avoiding wasted computation on irrecoverable states. This mechanism increases efficiency and prevents spurious rewards, ultimately leading to a more stable policy.
Figure 2. (a) Hand model from Meta XR Core SDK. (b) Objects: sphere, cube, and cylinder. (c) Sensors: GVS, LVS, closest and normal vector.
[…]
Illustration showing (a) the hand model with joint structures and colliders, (b) the three basic object types used for training—sphere, cube, and cylinder, and (c) the sensor setup including global voxel sensors (GVS), local voxel sensors (LVS), and closest point with normal vector visualization.
3.1. Hand Model and Objects
We utilize the Meta XR Core SDK hand model, which has 17 finger joints (23 total DoFs) and 19 non-self-colliding capsule colliders spread across the fingers and palm (Figure 2(a)). To learn various grip conditions, we use three primitive object shapes—sphere, cube, and cylinder (Figure 2(b))—each with a uniform mass of 0.01 kg. Since dropping an object results in early termination, the trained agent tends to apply excessive force to avoid dropping, even when this force does not match the target force. To prevent this mismatch, we mitigate the effect of mass by using lightweight objects.
Figure 3. Overview of our training pipeline. Rather than relying on motion-capture data, we generate diverse scenarios of varying object shapes, user inputs, and wrist motions; a three-phase curriculum progressively increases task complexity.
[…]
Curriculum learning framework diagram with three sequential stages: Finger Positioning, Intention Adaptation, and Dynamic Stabilization. Each stage introduces increasing complexity to train grip force control.
3.2. State and Action
@@
Hand model information $H=\{p,q,v,w,\alpha\}$ includes joint positions $p\in\mathbb{R}^{66}$, joint DoF angles $q\in\mathbb{R}^{23}$, joint linear velocities $v\in\mathbb{R}^{51}$, joint DoF velocities $w\in\mathbb{R}^{23}$, and joint DoF accelerations $\alpha\in\mathbb{R}^{23}$.
Object information $O=\{o_{g},v_{gvs},v_{lvs},c,n\}$ includes the gravity vector $o_{g}\in\mathbb{R}^{3}$ in wrist coordinates, a global voxel sensor $v_{gvs}\in\mathbb{R}^{450}$ attached to the wrist, employing 2 cm cells, local voxel sensors $v_{lvs}\in\mathbb{R}^{2160}$ attached to each collider, employing 0.5 cm cells, closest point vectors $c\in\mathbb{R}^{69}$ from relevant joints to the object’s surface, and corresponding surface normals $n\in\mathbb{R}^{69}$. Inspired by previous work (Zhang et al., 2021; Han et al., 2023), this dual-resolution sensing strategy provides a balance between detail and generality, minimizing overfitting to a specific object shape. Object information is illustrated in Figure 2.
@@ 3.3. Reward function
(3) $r_{p}=\sum_{j}{w^{j}\exp[-0.07||c_{t}^{j}||^{2}]}$
where $c_{t}^{j}$ is the distance vector from the $j$-th joint to the object surface, and $w^{j}$ is a weight emphasizing fingertip-first contact. We assign higher weights (0.0625) to the five end-effectors and their associated joints, and lower weights (0.03125) elsewhere. Specifically, joints or end-effectors within ~1cm of an object earn reward ¿0.9 before applying the weighted sum, encouraging natural contact.
@@
(4) $r=r_{f}+r_{p}$
Rather than ensuring all objects can always be lifted, our objective is to solely reflect user-intended force; insufficient force causes objects to slip. The user controls only the amount of force, not whether to grasp or release, making such slips a natural outcome by design.
4. Training
4.1. Training scenario generation
@@
4.2. Curriculum learning
Because these random, high-complexity scenarios pose significant training challenges, we devise a three-phase curriculum (Figure 3): Finger Positioning, Intention Adaptation, and Dynamic Stabilization. Each phase builds on the skills developed in the preceding phase, progressively training the agent to handle increasingly dynamic and realistic conditions.
4.2.1. Phase 1: Finger Positioning
This phase trains the agent to establish accurate finger contact. To minimize distractions, only object randomization is active; trigger variation and wrist movement are disabled. The wrist remains fixed with its pose randomly initialized, and user triggers are set to constant random values within [0,1]. This setup accelerates the agent’s mastery of foundational behaviors such as collision handling and initial grip formation.
4.2.2. Phase 2: Intention Adaptation
@@
4.2.3. Phase 3: Dynamic Stabilization